OpenStudy (anonymous):

A set P consists of all natural numbers between 1 and 100, inclusive. Use set-builder notation to define P.

OpenStudy (anonymous):

\[\left\{ x \in \mathbb{N} |1 \le x \le 100 \right\} \]

OpenStudy (anonymous):

Thanx

OpenStudy (anonymous):

You could also say: \[\left\{ x \in \mathbb{Z}^+ | x \le 100 \right\}\]